Table 5-7 1975 2005 $30$84 Mechanic s hourly wage Price Indez What has happened to a mechanics real hourly wage between 1975 and 20057 : 1) It rose by 180%. O 2) It fell by 18%. 3) It rose by 596. 4) It fell by about 64%. Save Previous Page Next Page Save All Responses Go to Submit Quiz 888A F2 F3 2 3 4 6
Answer #1

Real wage in 1975 = 30/1.2 = 25

Real wage in 2005 = 84/3.2 = 26.25

Increase in real wage = (26.25 - 25)/25 x 100 = 5%

So, answer is 3) It rose by 5%.